OpenStudy (anonymous):

what is the remainder when 2011 copies of '2011' (2011*2011...2011) Is divided by 7?

OpenStudy (ash2326):

Let's check for a small number first say if we take 4^4 and divide it by 7 Remainder when 4 is divided by 7=4 Remainder when 256(4^4) is divided by 7=4 so if take \(2011\times 2011\times 2011....2011\) and divide it by 7, we'll get the remainder same as when we divide 2011 by 7, remainder= remainder of 2011/7 if you find it out it comes as 2

OpenStudy (anonymous):

thanks!